# Better neural network expressivity: subdividing the simplex

### v1

@@
Abstract
This work studies the expressivity of $\operatorname{\mathsf{ReLU}}$ neural networks with a focus on their depth. A sequence of previous works showed that $\lceil\log_{2}(n+1)\rceil$ hidden layers are sufficient to compute all continuous piecewise linear ($\operatorname{\mathsf{CPWL}}$) functions on $\mathbb{R}^{n}$. Hertrich, Basu, Di Summa, and Skutella (NeurIPS ’21) conjectured that this result is optimal in the sense that there are $\operatorname{\mathsf{CPWL}}$ functions on $\mathbb{R}^{n}$, like the maximum function, that require this depth. We disprove the conjecture and show that $\lceil\log_{3}(n-1)\rceil+1$ hidden layers are sufficient to compute all $\operatorname{\mathsf{CPWL}}$ functions on $\mathbb{R}^{n}$.
@@ 1 Introduction
Functions represented by $\operatorname{\mathsf{ReLU}}$ networks are continuous and piecewise linear ($\operatorname{\mathsf{CPWL}}$), and every such function on $\mathbb{R}^{n}$ can be computed by a $\operatorname{\mathsf{ReLU}}$ network [2]. This naturally leads to complexity-theoretic questions, such as: what are the minimal size (i.e., number of neurons) and depth needed to represent specific $\operatorname{\mathsf{CPWL}}$ functions? Such questions have been extensively studied recently (see [2, 35, 19, 16, 33, 22, 20] and references within), but many major questions are still open, such as the minimum number of layers required to represent all $\operatorname{\mathsf{CPWL}}$ functions, which is the topic of this paper.
@@
Wang and Sun [34] showed that every function in $\operatorname{\mathsf{CPWL}}_{n}$ can be written as a linear combination of $\mathsf{MAX}_{n+1}$ functions applied to some affine functions: for every $f\in\operatorname{\mathsf{CPWL}}_{n}$, there are affine maps $A_{1},\ldots,A_{s}$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{n+1}$ and $\sigma_{1},\ldots,\sigma_{s}\in\{\pm 1\}$ so that
$f(x)=\sum_{i=1}^{s}\sigma_{i}\mathsf{MAX}_{n+1}(A_{i}(x)).$ (1)
As observed by Arora, Basu, Mianjy, and Mukherjee[2], this fact implies that
@@
Hertrich, Basu, Di Summa, and Skutella [19] conjectured that the $\lceil\log_{2}(n)\rceil$ upper bound on the required number of hidden layers for $\mathsf{MAX}_{n}$ is sharp. This conjecture led to a sequence of works that proved it in a variety of special cases. Mukherjee and Basu [28] are the first to prove the conjecture for the case $n=3$, i.e., showing that two hidden layers are necessary for $\mathsf{MAX}_{3}$. Haase, Hertrich, and Loho [16] proved the conjecture for general $n$ when the $\operatorname{\mathsf{ReLU}}$ network is allowed to use only integer weights. The techniques from [16] were extended by Averkov, Hojny, and Merkert [3], who proved logarithmic lower bounds for the case of fractional weights whose denominators do not have some fixed prime factor. The results of Valerdi [33] imply that the conjecture holds for the input convex neural networks ($\operatorname{\mathsf{ICNN}}$s) model, which was defined by Amos, Xu and Kolter [1] and studied in subsequent works (e.g. [7, 9, 10, 6, 20]). Stronger lower bounds for $\operatorname{\mathsf{ICNN}}$s and monotone $\operatorname{\mathsf{ReLU}}$ networks were shown by Bakaev, Brunck, Hertrich, Reichman, and Yehudayoff [4]. Grillo, Hertrich, and Loho proved an $\Omega(\log\log n)$ lower bound for $\operatorname{\mathsf{ReLU}}$ networks that “respect the singularities of $\mathsf{MAX}_{n}$” (see [14] for a formal definition). Their proof also implies the conjecture for $n=5$ for networks that “respect the singularities of $\mathsf{MAX}_{n}$”, a statement that was already shown in a computer-aided manner in [19].
Our Contribution.
@@
Compared to the previous upper bound [2], we improve the asymptotically required depth by a constant factor of $\ln(3)/\ln(2)\approx 0.63$. This surprising result shows that computing the maximum of $n$ numbers can be done in faster parallel time than $\log_{2}(n)$ even when all “max operations” have fan-in two. The summation operation allows to boost the parallel running time.
Naturally, our construction does not fall into any of the special cases under which the conjecture has been proven before. However, interestingly, as we only use weights with denominators that are a power of two, our construction falls into the setting of Averkov, Hojny, and Merkert [3], who proved a $\lceil\log_{3}(n)\rceil$ lower bound on the number of hidden layers required for $\mathsf{MAX}_{n}$ in the special case of $\operatorname{\mathsf{ReLU}}$ networks with weights that are decimal fractions. We almost match this lower bound (up to a difference of at most one layer).
It is noteworthy that Theorem 1 already improves the depth complexity of $\mathsf{MAX}_{5}$.
@@
This also implies that every function in $\operatorname{\mathsf{CPWL}}_{4}$ can be implemented with two hidden layers, sharpening the contrast to the $\operatorname{\mathsf{ICNN}}$ model, for which there are functions in $\operatorname{\mathsf{CPWL}}_{4}$ that need arbitrary large depth (see [33] for the proof using neighborly polytopes and [4] for a more detailed discussion).
@@
The proof of our result is short and elementary, we provide it in Section 2. However, it hides a beautiful geometric intuition that guided us in developing the proof and that might also be useful to further improve the current upper or lower bounds. We therefore explain this intuition in Section 3. The geometric intuition is based on relations between neural networks and polyhedral and tropical geometry, which were already studied in several previous works, see e.g., [36, 24, 25, 19, 22, 16, 5, 20, 33].
@@
We hope that our result initiates follow-up work on finding the true required depth to represent all $\operatorname{\mathsf{CPWL}}$ functions. There is a gap to close between $2$ and $\lceil\log_{3}(n-1)\rceil+1$ hidden layers. While we resolve the previously open question whether $\mathsf{MAX}_{5}$ can be represented with two hidden layers, it remains open whether the same is possible for $\mathsf{MAX}_{6}$.
Universal approximation and required size.
@@
Our paper is only concerned with the study of exact expressivity. The celebrated universal approximation theorems (see [11, 21] for initial versions and [23] for a version that encompasses $\operatorname{\mathsf{ReLU}}$) state that every continuous function on a bounded domain can be approximated with one hidden layer networks. A related work on approximating $\mathsf{MAX}_{n}$ is by Safran, Reichman, and Valiant [30]. Nevertheless, for a rigorous mathematical understanding of neural networks, it is inevitable to study the precise set of functions representable by different network architectures. Also, universal approximation theorems only work on bounded domains and might require a large number of neurons. In contrast, it is easy to verify that our construction needs only linearly many neurons to compute $\mathsf{MAX}_{n}$. It is, however, more tricky to determine the required size to represent an arbitrary $\operatorname{\mathsf{CPWL}}$ function with our construction, as this depends on the number $s$ of terms in the representation (1). The best known upper bound on $s$ was investigated in a series of works [17, 19, 8, 5] and is $s\in\mathcal{O}(p^{n+1})$, where $p$ is the number of different affine functions represented locally by the given $\operatorname{\mathsf{CPWL}}$ function. Another related stream of research is how one can trade depth and size against each other, with several results implying that mild increases in depth often allow significantly smaller neural networks [27, 32, 12, 2, 13, 29].
2 Direct proofs
2.1 Computing the maximum of five numbers with two hidden layers
@@
[…]
@@
With this representation of the maximum function, we can prove 3.
Proof of 3.
As each of the nine terms in the definition of $M$ can be computed with two hidden layers, the same is true for $M$ itself, and therefore, by 4, for $\mathsf{MAX}_{5}$. ∎
2.2 Improved upper bound on the depth
Computing $\mathsf{MAX}_{5}$ in $2$ hidden layers already allows us to compute $\mathsf{MAX}_{n}$ in $2\lceil\log_{5}(n)\rceil\approx 0.86\log_{2}(n)$ hidden layers. In fact, 4 can be pushed further. In this section, we show how to compute the maximum of $23^{n}+$ 2 numbers with $n+1$ hidden layers. This allows $\mathsf{MAX}_{n}$ to be computed in $\lceil\log_{3}(n-2)\rceil+1\approx 0.63\log_{2}(n)$ hidden layers.
@@ 2.2 Improved upper bound on the depth
Write $T_{a,b+4}=\mathsf{max}(x_{i+1},x_{i+2},x_{i+3},x_{i+4},T_{a,b})$, where $i=4a+b$. By Equation 2, this can be written as a linear combination of the nine functions $P_{1},P_{2},P_{3},P_{4},Q,R_{13},R_{14},R_{23}$ and $R_{24}$. Here, for example,
@@
$\mathcal{T}_{0,3^{n}+2}\subseteq\mathcal{T}_{1,3^{n}-3+2}\subseteq\mathcal{T}_{2,3^{n}-6+2}\subseteq\mathcal{T}_{3,3^{n}-9+2}\subseteq\cdots\subseteq\mathcal{T}_{3^{n-1},2}.$
@@
If the function $T_{0,3^{n-1}+2}$ can be computed in depth $k-1$, then $T_{3^{n-1},2}$ can be computed in depth $k$.
@@
Proof of Theorem 1.
The proof is by induction. The case $n=1$ is 3. For the induction step, we simply combine the previous two claims. Assume that $\mathsf{MAX}_{3^{n-1}+2}=T_{0,3^{n-1}+2}$ can be computed with $n$ hidden layers. By 7, we know that $T_{3^{n-1},2}$ can be computed with $n+1$ hidden layers. By 6, we now know that $T_{0,3^{n}+2}=\mathsf{MAX}_{3^{n}+2}$ can also be computed with $n+1$ hidden layers. ∎
Proof of Theorem 2.
This follows from Theorem 1, using the fact that every function in $\operatorname{\mathsf{CPWL}}_{n}$ can be written as a linear combination of maxima of $n+1$ affine terms [34]. ∎
@@
3 Geometric intuition
Even though it is not necessary for the proof of our main theorem, in this section we describe the geometric intuition behind our constructions. The geometric perspective may help to understand the constructions and may be useful in future research. The connection to geometry is based on a well-known fundamental equivalence between (positively homogeneous and convex) $\operatorname{\mathsf{CPWL}}$ functions and polytopes (see [36, 24, 25, 19, 22, 16, 5, 20, 33] and the references within), which we describe next.
3.1 Polytopes, support functions, and neural networks
A polytope is the convex hull of finitely many points in $\mathbb{R}^{n}$. We refer to [15] for standard terminology and theory of polytopes. Each polytope $P$ can be uniquely associated with its support function
@@ 3.1 Polytopes, support functions, and neural networks
The Newton polytope of $\mathsf{MAX}_{n}$ is the $(n-1)$-dimensional simplex $\Delta^{n-1}=\mathsf{conv}\{e_{1},\dots,e_{n}\}$ in $\mathbb{R}^{n}$, where $e_{i}$ is the $i$’th standard unit vector. Understanding the $\operatorname{\mathsf{ReLU}}$ depth complexity of $\mathsf{MAX}_{n}$ translates to a question about generating $\Delta^{n-1}$ in the following model (see e.g. [19, 33]). For two sets $P,Q\subseteq\mathbb{R}^{n}$, their Minkowski sum is defined as
@@
3.2 Upper bounds on the depth via subdivisions
By Lemma 8, we wish to write $\Delta^{n-1}$ as a formal Minkowski difference. We need to construct two polytopes $P,Q\in\operatorname{\mathcal{P}}_{k}$ for some small $k$, and show that $\Delta^{n-1}+P=Q$. This task is not so easy because $P$ and $Q$ should be somehow “coupled” for this equation to hold. We now explain how to replace the pair $(P,Q)$ by a subdivision of $\Delta^{n-1}$. The advantage is that subdivisions are easier to comprehend and hence to find.

### v2

@@
Abstract.
This work studies the expressivity of $\operatorname{\mathsf{ReLU}}$ neural networks with a focus on their depth. A sequence of previous works showed that $\lceil\log_{2}(n+1)\rceil$ hidden layers are sufficient to compute all continuous piecewise linear ($\operatorname{\mathsf{CPWL}}$) functions on $\mathbb{R}^{n}$. Hertrich, Basu, Di Summa, and Skutella (NeurIPS ’21 / SIDMA ’23) conjectured that this result is optimal in the sense that there are $\operatorname{\mathsf{CPWL}}$ functions on $\mathbb{R}^{n}$, like the maximum function, that require this depth. We disprove the conjecture and show that $\lceil\log_{3}(n-1)\rceil+1$ hidden layers are sufficient to compute all $\operatorname{\mathsf{CPWL}}$ functions on $\mathbb{R}^{n}$.
@@ 1. Introduction
Functions represented by $\operatorname{\mathsf{ReLU}}$ networks are continuous and piecewise linear ($\operatorname{\mathsf{CPWL}}$), and every such function on $\mathbb{R}^{n}$ can be computed by a $\operatorname{\mathsf{ReLU}}$ network [ABMM18]. This naturally leads to complexity-theoretic questions, such as: what are the minimal size (i.e., number of neurons) and depth needed to represent specific $\operatorname{\mathsf{CPWL}}$ functions? Such questions have been extensively studied recently (see [ABMM18, Wil18, HBDSS23, HHL23, Val24, HMST23, HL24] and references within), but many major questions are still open, such as the minimum number of layers required to represent all $\operatorname{\mathsf{CPWL}}$ functions, which is the topic of this paper.
@@
Wang and Sun [WS05] showed that every function in $\operatorname{\mathsf{CPWL}}_{n}$ can be written as a linear combination of $\mathsf{MAX}_{n+1}$ functions applied to some affine functions: for every $f\in\operatorname{\mathsf{CPWL}}_{n}$, there are affine maps $A_{1},\ldots,A_{s}$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{n+1}$ and $\sigma_{1},\ldots,\sigma_{s}\in\{\pm 1\}$ so that
(1) $f(x)=\sum_{i=1}^{s}\sigma_{i}\mathsf{MAX}_{n+1}(A_{i}(x)).$
As observed by Arora, Basu, Mianjy, and Mukherjee[ABMM18], this fact implies that
@@
Hertrich, Basu, Di Summa, and Skutella [HBDSS23] conjectured that the $\lceil\log_{2}(n)\rceil$ upper bound on the required number of hidden layers for $\mathsf{MAX}_{n}$ is sharp. This conjecture led to a sequence of works that proved it in a variety of special cases. Mukherjee and Basu [MB17] are the first to prove the conjecture for the case $n=3$, i.e., showing that two hidden layers are necessary for $\mathsf{MAX}_{3}$. Haase, Hertrich, and Loho [HHL23] proved the conjecture for general $n$ when the $\operatorname{\mathsf{ReLU}}$ network is allowed to use only integer weights. The techniques from [HHL23] were extended by Averkov, Hojny, and Merkert [AHM25], who proved logarithmic lower bounds for the case of fractional weights whose denominators do not have some fixed prime factor. The results of Valerdi [Val24] imply that the conjecture holds for the input convex neural networks ($\operatorname{\mathsf{ICNN}}$s) model, which was defined by Amos, Xu and Kolter [AXK17] and studied in subsequent works (e.g. [BSA+21, CSZ19, CSZ20, BSG+23, HL24]). Stronger lower bounds for $\operatorname{\mathsf{ICNN}}$s and monotone $\operatorname{\mathsf{ReLU}}$ networks were shown by Bakaev, Brunck, Hertrich, Reichman, and Yehudayoff [BBH+25]. Grillo, Hertrich, and Loho proved an $\Omega(\log\log n)$ lower bound for $\operatorname{\mathsf{ReLU}}$ networks that “respect the singularities of $\mathsf{MAX}_{n}$” (see [GHL25] for a formal definition). Their proof also implies the conjecture for $n=5$ for networks that “respect the singularities of $\mathsf{MAX}_{n}$”, a statement that was already shown in a computer-aided manner in [HBDSS23].
1.1. Our Contribution.
@@
Compared to the previous upper bound [ABMM18], we improve the asymptotically required depth by a constant factor of $\ln(2)/\ln(3)\approx 0.63$. This surprising result shows that computing the maximum of $n$ numbers can be done in faster parallel time than $\log_{2}(n)$ even when all “max operations” have fan-in two. The summation operation allows to boost the parallel running time.
Naturally, our construction does not fall into any of the special cases under which the conjecture has been proven before. However, interestingly, as we only use weights with denominators that are a power of two, our construction falls into the setting of Averkov, Hojny, and Merkert [AHM25], who proved a $\lceil\log_{3}(n)\rceil$ lower bound on the number of hidden layers required for $\mathsf{MAX}_{n}$ in the special case of $\operatorname{\mathsf{ReLU}}$ networks with weights that are decimal fractions. We almost match this lower bound (up to a difference of at most one layer).
It is noteworthy that Theorem˜1 already improves the depth complexity of $\mathsf{MAX}_{5}$.
@@
This also implies that every function in $\operatorname{\mathsf{CPWL}}_{4}$ can be implemented with two hidden layers, sharpening the contrast to the $\operatorname{\mathsf{ICNN}}$ model, for which there are functions in $\operatorname{\mathsf{CPWL}}_{4}$ that need arbitrary large depth (see [Val24] for the proof using neighborly polytopes and [BBH+25] for a more detailed discussion).
@@
The proof of our result is short and elementary, we provide it in Section˜2. However, it hides a beautiful geometric intuition that guided us in developing the proof and that might also be useful to further improve the current upper or lower bounds. We therefore explain this intuition in Section˜3. The geometric intuition is based on relations between neural networks and polyhedral and tropical geometry, which were already studied in several previous works, see e.g., [ZNL18, MS21, MCT21, HBDSS23, HMST23, HHL23, BGH25, HL24, Val24].
@@
We hope that our result initiates follow-up work on finding the true required depth to represent all $\operatorname{\mathsf{CPWL}}$ functions. There is a gap to close between $2$ and $\lceil\log_{3}(n-1)\rceil+1$ hidden layers. While we resolve the previously open question of whether $\mathsf{MAX}_{5}$ can be represented with two hidden layers, it intriguingly remains open whether the same is possible for $\mathsf{MAX}_{6}$.
We mention a similarity to the computational complexity of multiplying two $n\times n$ matrices. The trivial algorithm runs in time $O(n^{3})$. In his seminal work, Strassen showed that multiplying two $2\times 2$ matrices can be done with seven multiplications instead of eight [Str69]. This finite-dimensional improvement leads to a better algorithm with running time $O(n^{\log_{2}7})\approx O(n^{2.8})$. A long line of works (see [CU03, Cop82, GU18] and the many references within) found better and better upper bounds on the matrix multiplication exponent $\omega$ and the current best upper bound is $\omega\leq 2.371552$ [ADW+25]. Coming back to the maximum function, thinking of the depth parameter $d$ as fixed, we can trivially compute the maximum of $2^{d}$ numbers with $d$ hidden layers. The finite-dimensional ˜3 leads to a polynomial improvement. With $d$ hidden layers, we can compute the maximum of $\approx(5/2)^{d}$ numbers. With a more technical analysis this can be boosted to $\approx 3^{d}$. For the matrix multiplication exponent, we have the trivial lower bound $\omega\geq 2$. On the other hand, to the best of our knowledge it could be the case that two hidden layers suffice for computing the maximum of $n$ numbers for all $n$.
1.2. Universal approximation and required size.
@@
Our paper is concerned with the study of exact expressivity. The celebrated universal approximation theorems (see [Cyb89, HSW89] for initial versions and [LLPS93] for a version that encompasses $\operatorname{\mathsf{ReLU}}$) state that every continuous function on a bounded domain can be approximated with one hidden layer networks. Related to this, Safran, Reichman, and Valiant [SRV24] studied how to efficiently approximate $\mathsf{MAX}_{n}$. Nevertheless, for a rigorous mathematical understanding of neural networks, it is inevitable to study the precise set of functions representable by different network architectures. Also, universal approximation theorems only work on bounded domains and might require a large number of neurons. In contrast, it is easy to verify that our construction needs only linearly many neurons to compute $\mathsf{MAX}_{n}$. It is, however, more tricky to determine the required size to represent an arbitrary $\operatorname{\mathsf{CPWL}}$ function with our construction, as this depends on the number $s$ of terms in the representation (1). The best known upper bound on $s$ was investigated in a series of works [HLXZ20, HBDSS23, CGR22, BGH25] and is $s\in\mathcal{O}(p^{n+1})$, where $p$ is the number of different affine functions represented locally by the given $\operatorname{\mathsf{CPWL}}$ function. Another related stream of research is how one can trade depth and size against each other, with several results implying that mild increases in depth often allow significantly smaller neural networks [MPCB14, Tel16, ES16, ABMM18, EG24, SES22].
2. Direct proofs
2.1. Computing the maximum of five numbers with two hidden layers
@@
$\text{if }1\leftrightarrow 2,\text{ then }$ $P_{1}$ $\leftrightarrow\!\!\!\!$ $P_{2},\,$ $R_{13}$ $\leftrightarrow\!\!\!\!$ $R_{23},\,$ $R_{14}$ $\leftrightarrow\!\!\!\!$ $R_{24},$ other terms stay the same;
$\text{if }3\leftrightarrow 4,\text{ then }$ $P_{3}$ $\leftrightarrow\!\!\!\!$ $P_{4},\,$ $R_{13}$ $\leftrightarrow\!\!\!\!$ $R_{14},\,$ $R_{23}$ $\leftrightarrow\!\!\!\!$ $R_{24},$ other terms stay the same;
$\text{if }(1,2)\leftrightarrow(3,4),\text{ then }$ $P_{1}$ $\leftrightarrow\!\!\!\!$ $P_{3},\,$ $P_{2}$ $\leftrightarrow\!\!\!\!$ $P_{4},\,$ $R_{14}$ $\leftrightarrow\!\!\!\!$ $R_{23},$ other terms stay the same.
@@
With this representation of the maximum function, we can prove ˜3.
Proof of ˜3.
As each of the nine terms in the definition of $M$ can be computed with two hidden layers, the same is true for $M$ itself, and therefore, by ˜4, for $\mathsf{MAX}_{5}$. ∎
2.2. Improved upper bound on the depth
Computing $\mathsf{MAX}_{5}$ in $2$ hidden layers already allows us to compute $\mathsf{MAX}_{n}$ in $2\lceil\log_{5}(n)\rceil\approx 0.86\log_{2}(n)$ hidden layers. In fact, ˜4 can be pushed further. In this section, we show how to compute the maximum of $3^{n}+2$ numbers with $n+1$ hidden layers. This allows $\mathsf{MAX}_{n}$ to be computed in $\lceil\log_{3}(n-2)\rceil+1\approx 0.63\log_{2}(n)$ hidden layers.
@@ 2.2. Improved upper bound on the depth
Write $T_{a,b+4}=\mathsf{max}(x_{i+1},x_{i+2},x_{i+3},x_{i+4},T_{a,b})$, where $i=4a+b$. By Equation˜2, this can be written as a linear combination of the nine functions $P_{1},P_{2},P_{3},P_{4},Q,R_{13},R_{14},R_{23}$ and $R_{24}$. Here, for example,
@@
$\mathcal{T}_{0,3^{n}+2}\subseteq\mathcal{T}_{1,3^{n}-3+2}\subseteq\mathcal{T}_{2,3^{n}-6+2}\subseteq\mathcal{T}_{3,3^{n}-9+2}\subseteq\cdots\subseteq\mathcal{T}_{3^{n-1},2}.\qed$
@@
If the function $T_{0,3^{n-1}+2}$ can be computed in depth k1k-1, then $T_{3^{n-1},2}$ can be computed in depth $k$.
@@
Proof of Theorem˜1.
The proof is by induction. The case $n=1$ is ˜3. For the induction step, we simply combine the previous two claims. Assume that $\mathsf{MAX}_{3^{n-1}+2}=T_{0,3^{n-1}+2}$ can be computed with $n$ hidden layers. By ˜7, we know that $T_{3^{n-1},2}$ can be computed with $n+1$ hidden layers. By ˜6, we now know that $T_{0,3^{n}+2}=\mathsf{MAX}_{3^{n}+2}$ can also be computed with $n+1$ hidden layers. ∎
Proof of Theorem˜2.
This follows from Theorem˜1, using the fact that every function in $\operatorname{\mathsf{CPWL}}_{n}$ can be written as a linear combination of maxima of $n+1$ affine terms [WS05]. ∎
@@
3. Geometric intuition
Even though it is not necessary for the proof of our main theorem, in this section we describe the geometric intuition behind our constructions. The geometric perspective may help to understand the constructions and may be useful in future research. The connection to geometry is based on a well-known fundamental equivalence between (positively homogeneous and convex) $\operatorname{\mathsf{CPWL}}$ functions and polytopes (see [ZNL18, MS21, MCT21, HBDSS23, HMST23, HHL23, BGH25, HL24, Val24] and the references within), which we describe next.
3.1. Polytopes, support functions, and neural networks
A polytope is the convex hull of finitely many points in $\mathbb{R}^{n}$. We refer to [Grü03] for standard terminology and theory of polytopes. Each polytope $P$ can be uniquely associated with its support function
@@ 3.1. Polytopes, support functions, and neural networks
The Newton polytope of $\mathsf{MAX}_{n}$ is the (n1)(n-1)-dimensional simplex $\Delta^{n-1}=\mathsf{conv}\{e_{1},\dots,e_{n}\}$ in $\mathbb{R}^{n}$, where $e_{i}$ is the $i$’th standard unit vector. Understanding the $\operatorname{\mathsf{ReLU}}$ depth complexity of $\mathsf{MAX}_{n}$ translates to a question about generating $\Delta^{n-1}$ in the following model (see e.g. [HBDSS23, Val24]). For two sets $P,Q\subseteq\mathbb{R}^{n}$, their Minkowski sum is defined as
@@
3.2. Upper bounds on the depth via subdivisions
By Lemma˜8, we wish to write $\Delta^{n-1}$ as a formal Minkowski difference. We need to construct two polytopes $P,Q\in\operatorname{\mathcal{P}}_{k}$ for some small $k$, and show that $\Delta^{n-1}+P=Q$. This task is not so easy because $P$ and $Q$ should be somehow “coupled” for this equation to hold. We now explain how to replace the pair $(P,Q)$ by a subdivision of $\Delta^{n-1}$. The advantage is that subdivisions are easier to comprehend and hence to find.